### v1

@@ 5.1 CTQW on Complete, Cycle and Star Networks
To investigate the evolution of the CTQW on various network topologies, we first study the triangle network and compare its node occupation probabilities with the analytical results provided in Appendix A (see Fig. 9). We then examine the dynamics on simple networks such as the cycle, complete, and star topologies, as shown in Fig. 2. Star networks consists of two types of nodes: peripheral nodes, each with degree $1$, and a central hub node with degree $N-1$, so for this network, we consider two distinct initial conditions: in the first case, the walker is localized at the central hub node, and in the second case, it is initialized on one of the peripheral nodes. In contrast, both cycle and complete networks consist of nodes of uniform degree. Therefore, for these topologies, the CTQW is initially located on a randomly selected node due to the symmetry of the graph. In all the cases, the initially localized node refers to the node with probability 1 at the initial time step illustrated in Fig. 2.
@@
In the absence of noise, the CTQW on the cycle network as shown in Fig. 2(a), explores not only the initially localized node but also other nodes in the network. In contrast, for both the star and complete network topologies, the CTQW tends to remain localized near the initially occupied node, maintaining a high occupation probability at that node, as illustrated in Fig. 2(b)–(d). Moreover, the probability oscillations in the complete graph and the star network with hub node initialization exhibit a higher frequency compared to the star network with peripheral node initialization. For the star network with initial localization at a peripheral node in Fig. 3(c), the $\ell_{1}$-norm of coherence oscillates with lower frequency compared to the cases of the star with central node initialization and the complete graph, both of which show similar and more rapid oscillatory behavior as shown in Fig. 3(b),(d). This observation is consistent with previous findings by Xu et al. [48], who showed that transition probabilities are high-frequency revivals in the case of hub node excitation, along with strong long-term localization at the initial site for the star network. In addition to it, that study also demonstrated that the transport dynamics on a star graph with hub node excitation are equivalent to those on a complete graph of the same size. In contrast, for the cycle network, the $\ell_{1}$-norm of coherence displays fluctuations that do not go to zero, as shown in Fig. 3(a) and it exhibits a higher coherence, as measured by the $\ell_{1}$-norm of coherence, than both the complete and star networks. This can be attributed to the absence of high-degree nodes in the cycle network, which allows for delocalized dynamics and a higher value of coherence. These observations suggest that the dynamics of CTQW are highly sensitive to both the underlying network structure and the choice of initial localization.
Under Haken–Strobl noise, the occupation probabilities of the CTQW on the cycle graph as shown in Fig. 2(e) converge to a uniform distribution significantly faster than on the other topologies given in Fig. 2(f)-(h), indicating rapid delocalization and the onset of a steady state. This behavior resembles the equilibration of probabilities observed in classical continuous-time random walks [25]. The quantum coherence, quantified by the $\ell_{1}$-norm of coherence, also decays much more rapidly on the cycle topology as shown in Fig. 3(a) compared to the star and complete networks in Fig. 3(b)-(d). To determine the rate of coherence loss, we fitted the simulated coherence data using the stretched exponential Kohlrausch function [49]:
@@
Here, $\lambda$ is the effective decay rate that determines how quickly the function decreases(increases) over time, while $\beta$ is the stretching exponent that characterizes the nature of the decay. A value of $\beta=1$ corresponds to standard exponential decay, whereas $\beta<1$ coherence decay slower than exponential decay and for $\beta>1$ the quantum walker loses coherence faster than exponential decay. However, the overall behavior of the coherence decay is determined by the combined influence of both $\lambda$ and $\beta$. We fitted coherence data obtained for the Haken-Strobel and quantum stochastic decoherences as shown in Fig. 3(e),(f),(g),(h) and the extracted coefficients $\lambda$ and $\beta$, summarized in Table 1. These extracted coefficients $\lambda$ and $\beta$, further support our observations: the star network with the central node as the initial localization and the complete network exhibit slower coherence decay, indicating greater robustness against Haken–Strobl decoherence compared to both the cycle network and the star network initialized at a peripheral node. These findings conclude that, under Haken–Strobl noise, the star network initialized at a hub node and the complete network are more resilient in preserving quantum coherence than the cycle network and the star network initialized at the central node. Among all the network configurations considered, the cycle network performs the worst in maintaining quantum coherence for a long time and the star network with peripheral node initialization in a short time. However, it should be noted that for Haken-Strobl decoherence, all network topologies exhibit $\beta<1$ as in Table 1, indicating that the decay of coherence is slower than exponential in all cases.
Decoherence Network $\lambda$ $\beta$
Haken-Strobl Cycle 0.14 0.86
Star Hub 0.03 0.90
Star peripheral 0.10 0.71
Complete 0.03 0.90
Quantum Stochastic Cycle 0.08 2.03
Star Centre 0.28 2.62
Star peripheral 0.06 1.23
Complete 2.00 1.00
@@
Table 1: Coherence decay fitting parameters for cycle, star, and complete networks under Haken-Strobl and QSW.
@@
The evolution of CTQW under intrinsic decoherence is shown in Fig. 2(i)-(l), respectively, on cycle, star, and complete networks. From these figures, we observe distinct differences in the localization behavior, for instance, on star and complete topologies, CTQW tends to remain localized with a significantly high probability at the initial node compare than at other nodes in the network. In contrast, for the cycle topology, although the probabilities do not become exactly uniform, there is no strong localization on the initially localized node, unlike in star and complete topologies. Instead, the occupation probabilities of different nodes exhibit persistent oscillations over time (see Fig. 2(i)). As shown in Fig. 3(a)-(d) across all network types, coherence under intrinsic decoherence is preserved to varying extents. In the cycle network, the $\ell_{1}$-norm of coherence exhibits temporal fluctuations even at large times. For the star network with initial localization at a peripheral node, coherence exhibits damped oscillations before settling into a steady state value. Interestingly, for both the complete network and the star network with central hub node initialization, the coherence also converges to the same steady state value observed in the peripheral initialization case. Although the cycle network retains some coherence under intrinsic decoherence, the level of preservation is lower compared to the star and complete topologies. Notably, the cycle exhibits oscillatory behavior in both probability and coherence dynamics under intrinsic decoherence, a feature that is absent in the star and complete networks.
@@
Figure 3: $\ell_{1}$-norm of coherence with time for simple topologies with network size N = 10. In the first row, $\ell_{1}$-norm of coherence is plotted for all types of decoherence for cycle (a), star network with hub node initialization (b) and peripheral node initilization (c), and complete network (d). The second row presents the coherence decay profiles for both Haken–Strobl noise and QSW, with the data fitted using the proposed decay model in Eq. (11), with the best-fit parameters summarized in Table 1.
@@
In the case of QSW decoherence, the decay of coherence is faster for complete network than other topologies. However, decay of coherence under QSW decoherence is much faster than Haken-Strobl for all network topologies, as shown in Table 1. Star graph with peripheral node initialization takes longer to reach the same probability distribution and exhibits a slower decay of coherence compared to other topologies and the star graph with center initialization, as seen in Fig. 2(m)-(p) and Fig. 3(e)-(h). This observation is further supported by exponents $\lambda$ and $\beta$ as shown in Table 1, which shows that the coherence takes a longer time to decay in the star topology than cycle and complete topologies, indicating greater stability of this topology under QSW induced decoherence.
@@
The behavior of quantum-classical distance, fidelity with the initial state, and von Neumann entropy is presented in Appendix B for networks of size $N=50$ in Fig. 12 and Fig. 13, and for networks of size $N=10$ in Fig. 10 and Fig. 11. The $N=50$ results are particularly informative, clearly distinguishing the relative stability of different topologies. Under Haken-Strobl noise and intrinsic decoherence, the star network (with hub node initialization) and the complete graph exhibit high stability, indicated by elevated quantum-classical distances and fidelities, along with low entropy. For the QSW, the star network initialized at a peripheral node shows improved stability, as seen from its high fidelity and low entropy. Interestingly, the quantum-classical distance highlights the cycle network as especially robust against classicalization.
5.2 CTQW on Complex Network Topologies
@@
To further validate our observations, we extend our analysis to more complex topologies to investigate how different network structures influence the stability of quantum systems. Specifically, we study the evolution of CTQW on Erdős–Rényi, small-world, and scale-free networks. As the degree of the initially localized node can significantly affect the dynamics, we choose a high-degree node as the initial localization in the heterogeneous scale-free network. In contrast, the Erdős–Rényi and small-world networks have nodes with approximately similar average degrees, so the impact of the initial node selection is expected to be comparatively minimal in these cases. In all cases, the initially localized node is defined as the node with probability $1$ at the initial time step (see Fig. 4).
@@
Figure 5: Time evolution of the $\ell_{1}$-norm of coherence for complex network topologies with size $N=100$. Coherence decay under Haken–Strobl noise and quantum stochastic walk (QSW) is modeled using the decay function given in Eq. (11), with best-fit parameters summarized in Table 2. The $\ell_{1}$-norm of coherence is plotted for the Erdős–Rényi network (a), small-world network (b), and scale-free network (c) under four conditions: no noise (blue), Haken–Strobl noise (red), intrinsic decoherence (cyan), and quantum stochastic walk (purple). These curves illustrate how coherence either diminishes or persists over time, depending on network structure and decoherence model.
@@
Fig. 4 shows that the scale-free network exhibits distinct behavior compared to the Erdős–Rényi and small-world networks, particularly in the noiseless case, as illustrated in Fig. 4(a),(e),(i). Node centralities in scale-free networks vary significantly due to their inherently heterogeneous structure. As a result, the evolution of CTQW on such networks depends strongly on the centrality of the node where the walker is initially localized. For instance, when the initially localized node has the highest degree but not the highest closeness centrality, the occupation probabilities at the highest-degree and highest-closeness nodes exhibit an anti-phase synchronization pattern (see Appendix B Fig. 14). The probability amplitudes of anti-phase synchronized nodes vary depending on the specific configuration of the network structure. Notably, strong localization at the initially occupied hub node is observed in heterogeneous networks, such as, the scale-free network Fig. 4(i) and star networks. This behavior is consistent with the previous study by Xu et al. [48] who showed that, in the thermodynamic limit ($\lim{N\to\infty}$), transition probabilities in star networks tend to remain localized at the initially localised site. In contrast, Erdős–Rényi and small-world networks lack pronounced degree heterogeneity and central hubs, leading to a more delocalized evolution of the CTQW. Interestingly, however, Fig. 5 shows that, in the absence of decoherence, both small-world and Erdős–Rényi networks maintain a higher $\ell_{1}$-norm of coherence compared to the scale-free network. This highlights the influence of network structure, suggesting that while hubs facilitate localization, they may simultaneously contribute to reduced coherence in noise-free conditions.
@@
Under the influence of Haken–Strobl noise, the occupation probability at the initially localized node decays rapidly in both the Erdős–Rényi and small-world networks Fig. 4(b),(f), eventually leading to a uniform distribution. This renders the initially localized node indistinguishable from others, indicating complete delocalization across the network. In contrast, the scale-free network does not exhibit complete delocalization Fig. 4(j); although the occupation probability at the initially localized node decays over time, the distribution remains uneven, and the node probabilities do not converge to a uniform state.
Furthermore, we extracted the coherence decay coefficients $\lambda$ and $\beta$ (summarized in Table 2) by fitting the coherence data as shown in the red curves in Fig. 5 using Eq. (11). The red curves in Fig. 5(c) show that the scale-free network preserves quantum coherence for a longer duration under Haken–Strobl noise, as indicated by its relatively slow decay, indicated by a low $\beta$ value.
In the presence of intrinsic decoherence, the CTQW exhibits partial localization at the initially localized node in the scale-free network Fig. 4(k), similar to the behavior observed in the noiseless case, but with a fixed probability. Localization behavior of CTQW is not observed in the Erdős–Rényi and small-world networks, as illustrated in Fig. 4(c),(g). Interestingly, the $\ell_{1}$-norm of coherence is not completely destroyed under intrinsic decoherence for any of the network topologies. This suggests that, despite inducing partial localization in the scale-free network, intrinsic decoherence preserves quantum coherence across all topologies, ensuring stability for all topologies.
@@
On the other hand, QSW dynamics lead to a complete delocalization of the quantum walker across all network types. As shown in Fig. 4(d),(h),(l), node probabilities tend to equalize, erasing the initial localization, particularly evident in the scale-free network, where the initial localization is entirely lost. Coherence analysis of $\lambda$ and $\beta$ from Table 2, and purple curves in Fig. 5 further shows that QSW induces the fastest coherence decay among all the decohernece. Although Erdős–Rényi networks retain more coherence compared to other topologies under QSW, overall coherence degrades rapidly across all complex network topologies. However, this decay is not faster than exponential, as indicated by $\beta<1$.
Decoherence Network $\lambda$ $\beta$
Haken-Strobl Erdős-Rényi 0.10 0.84
Small-World 0.17 0.96
Scale-Free 0.24 0.37
Quantum Stochastic Erdős-Rényi 0.71 0.45
Small-World 1.53 0.90
Scale-Free 0.96 0.62
@@
Table 2: Coherence decay fitting parameters for Erdős-Rényi, small-world and scale-free under Haken-Strobl and QSW
@@
Further to assess the system’s stability, we utilize a set of standard metrics from open quantum system analysis, such as quantum-classical distance, von Neumann entropy, and fidelity. The quantum-classical distance in Fig. 6(a)-(d) tells that the scale-free network exhibits greater resistance to classicalization under decoherence compared to the Erdős–Rényi and small world network, since as time progresses, scale-free has more quantum classical distance than other topologies underlying all decoherence dynamics. It maintains a significantly higher quantum-classical distance in the presence of both QSW and Haken-strobl both of which tend to classicalize the system over long timescales. For intrinsic decoherence in Fig. 6(c), the distance does not decay to zero, indicating that classicalization does not occur and quantum features are consistently preserved for all networks. For the noiseless scenario, across all network topologies, the quantum-classical distance converges to the theoretical value $\lim_{t\to\infty}D_{QC}(t)=1-\frac{1}{N}$ [37], as represented by the dashed line in Fig. 6(a).
@@
The fidelity plots in Fig. 6(e)-(h) show that the scale-free network topology exhibits localization near the initial state since it has high fidelity with the initial state. Higher the fidelity lesser the evolved state has deviated from the initial state. On contrast, the Erdős-Rényi and small-world networks have a little less fidelity across various decoherences.
@@
Von Neumann entropy plots Fig. 7(a)-(c) show that entropy is high for Erdős-Rényi in all decoherence. Entropy is least for scale-free in the case of Haken-Strobl and QSW, showing scale-free is not much deviated from the pure state as compared to Erdős-Rényi. This implies that scale-free is more resilient to all noises and Erds-Rényi is more prone to mixed states than other topologies. For the noiseless case, the entropy always remains zero.
6 Conclusion

### v2

@@ V.1 CTQW on Complete, Cycle and Star Networks
To study the evolution of the CTQW on various network topologies, we first examine the dynamics on simple networks such as the cycle, complete, and star topologies. Star networks consists of two types of nodes: peripheral nodes, each with degree $1$, and a central hub node with degree N1N-1, so for this network, we consider two distinct initial conditions: in the first case, the walker is localized at the central hub node, and in the second case, it is initialized on one of the peripheral nodes. In contrast, both cycle and complete networks consist of nodes of uniform degree. Therefore, for these topologies, the CTQW is initially located on a randomly selected node due to the symmetry of the graph. In all the cases, the initially localized node refers to the node with probability 1 at the initial time step.
To probe the system’s stability, we first conduct a detailed scaling analysis of the star network with central node initialization, studying how its $\ell_{1}$-norm of coherence depends on both network size N and decoherence rate. As shown in Fig. 2(a), for Haken-Strobl noise the time taken for coherence decay increases with network size, showing that stability improves as $N$ increases. For QSW, however, the opposite is true. Fig. 2(b) shows that as $N$ increases, the coherence decays faster, indicating that larger networks are little less stable under this decoherence. For intrinsic decoherence, Fig. 2(c) shows that as $N$ increases, the coherence saturates faster and to a higher steady-state coherence value, suggesting greater stability. Although smaller networks exhibit a slight peak before stabilizing, the reduced overall decay for larger $N$ shows that stability is enhanced. Next, we study the dependence on decoherence rates for a fixed network size of $N=10$. As shown in Fig. 2(d) for Haken-Strobl noise and Fig. 2(e) for QSW, stability uniformly decreases as the respective rates $\gamma$ and $p$ increase. For intrinsic decoherence Fig. 2(f) shows that while larger $\gamma$ value dampen initial oscillations and lead to faster saturation, the final steady-state coherence value is dependent on the network size, not on the decoherence rate. Having studied this baseline scaling, we now proceed to a detailed comparison of the dynamics across the complete, cycle, star topologies with fixed size $N=10$. For this and all subsequent simulations, fixed decoherence rates of $\gamma=0.1$ and $p=0.1$, as this scaling analysis confirms they are large enough to make the decoherence effects clearly visible.
@@
In the absence of noise, the CTQW on the cycle network as shown in Fig. 3(a), explores not only the initially localized node but also other nodes in the network. In contrast, for both the star and complete network topologies, the CTQW tends to remain localized near the initially occupied node, maintaining a high occupation probability at that node, as illustrated in Fig. 3(b),(d). As shown by Razzoli et al. [razzoli2022universality], the dynamics of a CTQW initialized on a fully connected hub (a node with degree N-1) are universal, meaning they are determined only by the size of the network, N, and are independent of the rest of the graph’s structure. Our simulations shows that this universality is not only true for noisless case, but also true for Haken-Strobl and intrinsic decoherence as shown in Fig. 3(f),(h),(j),(l). Moreover, the probability oscillations in the complete graph and the star network with hub node initialization are identical and exhibit a higher frequency compared to the star network with peripheral node initialization. For the star network with initial localization at a peripheral node in Fig. 4(c), the $\ell_{1}$-norm of coherence oscillates with lower frequency compared to the cases of the star with central node initialization and the complete graph, both of which show similar and more rapid oscillatory behavior as shown in Fig. 4(b),(d). This observation is consistent with previous findings by Xu et al. [xu2009exact], who showed that transition probabilities are high-frequency revivals in the case of hub node excitation, along with strong long-term localization at the initial site for the star network. In addition to it, that study also demonstrated that the transport dynamics on a star graph with hub node excitation are equivalent to those on a complete graph of the same size. In contrast, for the cycle network, the $\ell_{1}$-norm of coherence displays fluctuations that do not go to zero, as shown in Fig. 4(a) and it exhibits a higher coherence, as measured by the $\ell_{1}$-norm of coherence, than both the complete and star networks. This can be attributed to the absence of high-degree nodes in the cycle network, which allows for delocalized dynamics and a higher value of coherence. These results suggest that the dynamics of CTQW are highly sensitive to both the underlying network structure and the choice of initial localization. Under Haken–Strobl noise, the occupation probabilities of the CTQW on the cycle graph as shown in Fig. 3(e) converge to a uniform distribution significantly faster than on the other topologies given in Fig. 3(f)-(h), indicating rapid delocalization and the onset of a steady state. This behavior resembles the equilibration of probabilities observed in classical continuous-time random walks [mulken2011continuous]. The system’s convergence to the maximally mixed state $\frac{\mathbb{I}}{N}$ (equilibration of probabilities and $\ell_{1}$-norm of coherence decays to zero) is consistent with the fact that both the Haken-Strobl noise and QSW Liouvillians are unital and Davies irreducible [davies1976quantum] for connected graphs. As detailed in Appendix  A.1 and  A.2, we verify irreducibility using the algebraic criteria of Zhang and Barthel [zhang2024criteria], which guarantees a unique and maximally mixed steady state. The quantum coherence, quantified by the $\ell_{1}$-norm of coherence, also decays much more rapidly on the cycle topology as shown in Fig. 4(a) compared to the star and complete networks in Fig. 4(b)-(d). To determine the rate of coherence loss, we fitted the simulated coherence data using the stretched exponential Kohlrausch function [kohlrausch1854theorie]:
@@
Here, $\lambda$ is the effective decay rate that determines how quickly the function decreases (increases) over time, while $\beta$ is the stretching exponent that characterizes the nature of the decay. A value of $\beta=1$ corresponds to standard exponential decay, whereas $\beta<1$ coherence decay slower than exponential decay and for $\beta>1$ the quantum walker loses coherence faster than exponential decay. However, the overall behavior of the coherence decay is determined by the combined influence of both $\lambda$ and $\beta$, which can be calculated as a single characteristic relaxation time, $\tau$, which represents the mean lifetime of the coherence from its normalized initial value. given by  [gueguen2015relationship]:
$\tau=\frac{1}{\beta\lambda}\,\Gamma\!\left(\frac{1}{\beta}\right)$ (12)
@@
We fitted coherence data obtained for the Haken-Strobl noise and QSW as shown in Fig. 4(e),(f),(g),(h) and the extracted coefficients $\lambda$, $\beta$ and their characteristic relaxation time, summarized in Table 1. These characteristic relaxation time $\tau$, further support our observations: the star network with the central node as the initial localization and the complete network exhibit slower coherence decay characterized by higher and identical $\tau$, indicating greater robustness against Haken–Strobl decoherence compared to both the cycle network and the star network initialized at a peripheral node. Moreover this identical $\lambda$ and $\beta$ as shown in Table 1 for the fully connected, initially localized node, show that the universality discussed earlier holds for Haken-strobl noise, applying not only to probability but to coherence as well. Among all the network configurations considered, the cycle network performs the worst in maintaining quantum coherence for a long time and the star network with peripheral node initialization in a short time. However, it should be noted that for Haken-Strobl decoherence, all network topologies exhibit $\beta<1$ as in Table 1, indicating that the decay of coherence is slower than exponential in all cases.
Decoherence Network $\lambda$ $\beta$ $\tau$
Haken-Strobl Cycle 0.14 0.86 7.72
Star Hub 0.03 0.90 35.07
Star peripheral 0.10 0.71 12.49
Complete 0.03 0.90 35.07
Quantum Stochastic Cycle 0.08 2.03 11.08
Star Centre 0.28 2.62 3.17
Star peripheral 0.06 1.23 15.58
Complete 2.00 1.00 0.50
@@
Table 1: Coherence decay fitting parameters for cycle, star, and complete networks under Haken-Strobl and QSW decoherence.
@@
The evolution of CTQW under intrinsic decoherence is shown in Fig. 3(i),(j),(k),(l), respectively, on cycle, star, and complete networks. From these figures, we observe distinct differences in the localization behavior, for instance, on star and complete topologies, CTQW tends to remain localized with a significantly high probability at the initial node compare than at other nodes in the network. In contrast, for the cycle topology, although the probabilities do not become exactly uniform, there is no strong localization on the initially localized node, unlike in star and complete topologies. Instead, the occupation probabilities of different nodes exhibit persistent oscillations over time (see Fig. 3(i)). As shown in Fig. 4(a)-(d) across all network types, coherence under intrinsic decoherence is preserved to varying extents. In the cycle network, the $\ell_{1}$-norm of coherence exhibits temporal fluctuations even at large times. For the star network with initial localization at a peripheral node, coherence exhibits damped oscillations before settling into a steady state value. Interestingly, for both the complete network and the star network with central hub node initialization, the coherence also converges to the same steady state value observed in the peripheral initialization case. Although the cycle network retains some coherence under intrinsic decoherence, the level of preservation is lower compared to the star and complete topologies. Notably, the cycle exhibits oscillatory behavior in both probability and coherence dynamics under intrinsic decoherence, a feature that is absent in the star and complete networks.
@@
Figure 4: $\ell_{1}$-norm of coherence with time for simple topologies with network size N = 10. In the first row, $\ell_{1}$-norm of coherence is plotted for all types of decoherence for cycle (a), star network with hub node initialization (b) and peripheral node initialization (c), and complete network (d). The second row presents (e)-(f) the coherence decay profiles for both Haken–Strobl noise and QSW, with the data fitted using the proposed decay model in Eq. (11), with the best-fit parameters summarized in Table 1.
@@
In the case of QSW decoherence, the decay of coherence is faster for complete network than other topologies. However, decay of coherence under QSW decoherence is much faster than Haken-Strobl for all network topologies, as shown in Table 1 and the stability ranking between topologies is also inverted. Star graph with peripheral node initialization takes longer to reach the same probability distribution and exhibits a slower decay of coherence compared to other topologies and the star graph with center initialization, as seen in Fig. 3(m)-(p) and Fig. 4(e)-(h). This observation is further supported by $\tau$ as shown in Table 1, which shows that the coherence takes a longer time to decay in the star topology with initialization on peripheral node than all the other topologies, indicating greater stability of this topology under QSW induced decoherence. This complete reversal suggests that stability under QSW is inversely related to node connectivity. Because QSW is a link-based decoherence model (Eq. (4)), it appears to more dissipation as the connections in the network increases, explaining why the highly-connected complete graph and star-hub configurations are so fragile, while the minimally connected peripheral node is the most stable.
@@
The behavior of quantum-classical distance, fidelity with the initial state, and von Neumann entropy is presented in Appendix B for networks of size $N=10$ Fig. 9 and Fig. 10. Under Haken-Strobl noise and intrinsic decoherence, the star network (with hub node initialization) and the complete graph exhibit high stability, indicated by elevated quantum-classical distance and fidelity, along with low entropy for Haken-Strobl. For the QSW, the star network initialized at a peripheral node shows improved stability, as seen from its high fidelity and low entropy. Interestingly, the quantum-classical distance highlights the cycle network as especially robust against classicalization.
V.2 CTQW on Complex Network Topologies
@@
We extend our analysis to more complex topologies to investigate how different network structures influence the stability of quantum systems. Specifically, we study the evolution of CTQW on Erdős–Rényi, small-world, and scale-free networks. As the degree of the initially localized node can significantly affect the dynamics (see Appendix C Fig. 11), we choose a high-degree node as the initial localization in the heterogeneous scale-free network. In contrast, the Erdős–Rényi and small-world networks have nodes with approximately similar average degrees, so the impact of the initial node selection is expected to be comparatively minimal in these cases. In all cases, the initially localized node is defined as the node with probability $1$ at the initial time step (see Fig. 5).
Erdős–Rényi Network
Refer to captionProbability
Small-World Network
Refer to captionProbability
Scale-Free Network
Refer to captionProbability
@@
Figure 6: Time evolution of the $\ell_{1}$-norm of coherence for complex network topologies with size N = 100. The coherence decay under Haken–Strobl noise and QSW is modeled using the decay function given in Eq. (11), with the best-fit parameters summarized in Table 2. The $\ell_{1}$-norm of coherence is plotted for the Erdős–Rényi network (a), small-world network (b), and scale-free network (c) under four conditions: no noise (blue), Haken–Strobl noise (red), intrinsic decoherence (cyan), and quantum stochastic walk (purple). These curves illustrate how coherence diminishes or preserve over time depending on the network structure and the specific decoherence model.
@@
Fig. 5 shows that the scale-free network exhibits distinct behavior compared to the Erdős–Rényi and small-world networks, particularly in the noiseless case, as illustrated in Fig. 5(a),(e),(i). Node centralities in scale-free networks vary significantly due to their inherently heterogeneous structure. As a result, the evolution of CTQW on such networks depends strongly on the centrality of the node where the walker is initially localized. For instance, when the initially localized node has the highest degree but not the highest closeness centrality, the occupation probabilities at the highest-degree and highest-closeness nodes exhibit an anti-phase synchronization pattern (see Appendix C Fig. 11). The probability amplitudes of anti-phase synchronized nodes vary depending on the specific configuration of the network structure. Notably, strong localization at the initially occupied hub node is observed in heterogeneous networks, such as, the scale-free network Fig. 5(i) and star networks. This behavior is consistent with the previous study by Xu et al. [xu2009exact] who showed that, in the thermodynamic limit ($\lim{N\to\infty}$), transition probabilities in star networks tend to remain localized at the initially localized site. In contrast, Erdős–Rényi and small-world networks lack pronounced degree heterogeneity and central hubs, leading to a more delocalized evolution of the CTQW. Interestingly, however, Fig. 6 shows that, in the absence of decoherence, both small-world and Erdős–Rényi networks maintain a higher $\ell_{1}$-norm of coherence compared to the scale-free network. This highlights the influence of network structure, suggesting that while hubs facilitate localization, they may simultaneously contribute to reduced coherence in noiseless condition.
@@
Under the influence of Haken–Strobl noise, the occupation probability at the initially localized node decays rapidly in both the Erdős–Rényi and small-world networks Fig. 5(b),(f), eventually leading to a uniform distribution. This renders the initially localized node indistinguishable from others, indicating complete delocalization across the network. In contrast, the scale-free network does not exhibit complete delocalization in short time Fig. 5(j). Although the occupation probability at the initially localized node decays over time, the distribution remains uneven and converges toward uniformity at a significantly slower rate than that observed in the Erdős–Rényi and small-world networks.
@@
Furthermore, we extracted the coherence decay coefficients $\lambda$ and $\beta$ by fitting the coherence data as shown in the red curves in Fig. 6 using Eq. (11). From these, we then calculated the relaxation time, $\tau$ using Eq. 12, and all three parameters ($\lambda$$\beta$ and $\tau$) are summarized in Table 2. The red curves in Fig. 6(c) show that the scale-free network preserves quantum coherence for a longer duration under Haken–Strobl noise, as indicated by its relatively slow decay, showed by a high relaxation time. In the presence of intrinsic decoherence, the CTQW exhibits partial localization at the initially localized node in the scale-free network Fig. 5(k), similar to the behavior observed in the noiseless case, but with a fixed probability. Localization behavior of CTQW is not observed in the Erdős–Rényi and small-world networks, as illustrated in Fig. 5(c),(g). Interestingly, the $\ell_{1}$-norm of coherence is preserved under intrinsic decoherence for any of the network topologies. This suggests that, despite inducing partial localization in the scale-free network, intrinsic decoherence preserves quantum coherence across all topologies, ensuring stability for all topologies. On the other hand, QSW decoherence lead to a complete delocalization of the quantum walker across all network types. As shown in Fig. 5(d),(h),(l), node probabilities tend to equalize, erasing the initial localization, particularly evident in the scale-free network, where the initial localization is entirely lost. Relaxation time ($\tau$) from Table 2, and purple curves in Fig. 6 further shows that QSW induces the fastest coherence decay among all the decoherence. Although Erdős–Rényi networks retain coherence for more time compared to other topologies under QSW, overall coherence degrades rapidly across all complex network topologies. However, this decay is not faster than exponential as in simple topologies discussed in Sec. V.1, as indicated by $\beta<1$.
Decoherence Network $\lambda$ $\beta$ $\tau$
Haken-Strobl Erdős-Rényi 0.10 0.84 10.96
Small-World 0.17 0.96 5.99
Scale-Free 0.24 0.37 17.43
Quantum Stochastic Erdős-Rényi 0.71 0.45 3.49
Small-World 1.53 0.90 0.69
Scale-Free 0.96 0.62 1.50
@@
Table 2: Coherence decay fitting parameters for Erdős-Rényi, small-world and scale-free under Haken-Strobl and QSW decoherence.
@@
Further to assess the system’s stability, we utilize a set of standard metrics, quantum-classical distance, von Neumann entropy, and fidelity. The quantum-classical distance in Fig. 9(a)-(d) tells that the scale-free network exhibits greater resistance to classicalization under decoherence compared to the Erdős–Rényi and small world network, since as time progresses, scale-free has higher quantum classical distance than other topologies underlying all decoherence dynamics. It shows a slower decay of quantum-classical distance in the presence of both QSW and Haken-strobl both of which classicalizes the system over long timescales. For intrinsic decoherence in Fig. 9(c), the distance does not decay to zero, indicating that classicalization does not occur and quantum features are consistently preserved for all networks. For the noiseless scenario, across all network topologies, the quantum-classical distance converges to the theoretical value $\lim_{t\to\infty}D_{QC}(t)=1-\frac{1}{N}$ [gualtieri2020quantum], as represented by the dashed line in Fig. 9(a).
@@
The fidelity plots in Fig. 9(e)-(h) show that the scale-free network topology exhibits localization near the initial state since it has high fidelity with the initial state. Higher the fidelity lesser the evolved state has deviated from the initial state. On contrast, the Erdős-Rényi and small-world networks have a little less fidelity across various decoherence.
@@
Von Neumann entropy plots Fig. 10(a)-(c) show that entropy is high for Erdős-Rényi in all decoherence. Entropy is least for scale-free in the case of Haken-Strobl and QSW, showing scale-free is not much deviated from the pure state as compared to Erdős-Rényi. This implies that scale-free is more resilient to all noises and Erdős-Rényi is more prone to mixed states than other topologies. For the noiseless case, the entropy always remains zero.
VI Conclusion